Identify the expression that completes the equation so that it is an identity.
cos x/1+sin x = ?

Thanks!

Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

Identify the expression that completes the equation so that it is an identity.
cos x/1+sin x = cos(x) + sin(x) = cos(x) + sqrt(1 - sin^2(x))

Answer by Edwin McCravy(20056)